The length of a rectangular garden is 6 more than the width. The area is 27 sq. units. Find the dimensions of the garden using the quadratic formula.
Found 2 solutions by Cromlix, solver91311:
Answer by Cromlix(4381)

Width = x
Length = x + 6
x(x + 6) = 27
x^2 + 6x - 27 = 0
(x + 9)(x - 3) = 0
x = -9 (no answer)
x = 3
Length = x + 6
= 3 + 6
= 9 units
Width = x
= 3 units
Area = 9 units x 3 units = 27 units^2
Hope this helps
:-)

Answer by solver91311(24713)



Let represent the width. Then represents the length. Area is length times width, so which is to say .

Use the quadratic formula:




where , , and

Throw out the absurd negative root and then calculate

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism